vanscpy bank pays 7.8% simple interest on it's investment account.if case bank pays interest on it's accounts compounded annually,what rate should the bank set if it wants to match vancoy bank over an investment horizon of eight years
Answer #1

Let the rate compounded annually be x

Future value under both cases will be same after 8 years

Amount + Amount*7.8%*8 = Amount*(1+x)^8

Amount(1+0.624)= Amount*(1+x)^8

1.624 = (1+x)^8

1+x = 1.062489

Hence, x = 6.2489%

i.e. 6.25%

Hence, required rate compounded annually = 6.25%